FIND THE MISSING CORDINATES 2X+4Y=24 (8, ) ( ,-2)
Answer by bsmath(21)

The x value is 8, so 2*8+4*y=24. So 4*y=24-16, 4*y=8, y=2. For the next one, the y value is -2,
so 2*x+4*-2=24, 2*x=24+8, 2*x=32, x=16.